OpenStudy (anonymous):

Find the angle between Oy axis and vector a{1,squareroot of 2,1} ? Thanks!

OpenStudy (anonymous):

Angle between them are \[\theta=\cos^{-1}\frac{\sqrt2}{2}=\frac{\pi}{4}\]

OpenStudy (anonymous):

Nice! How do you count that?

OpenStudy (anonymous):

Find unit vector along the a<1,sqrt(2),1> which is 0.5<1,sqrt(2),1> take dot product with unit vector along OY